Please help me in this question i got it for h.w and i need it for tomorrow please help! Thanks so much i really appreciate it!
Question: Why is it impossible for any point to lie on both y=3x+5 and on y=3x-2?
Please help and show how you got to it. Thanks! :)

Found 2 solutions by jim_thompson5910, scott8148:
Answer by jim_thompson5910(35256)

If y=3x+5 and y=3x-2, then 3x+5 = 3x-2 ( I started with y=3x-2, then replaced 'y' with '3x+5' on the left side)

Solve this equation to get


3x+5=3x-2


3x=3x-2-5


3x-3x=-2-5


0x=-2-5


0x=-7


0=-7


Since this equation is NEVER true for any x value, this means that there are no solutions.


So because there are no solutions for x, there are no solutions for y as well. Therefore, there are no points that lie on both lines at the same time.


So the two lines never intersect.

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

Answer by scott8148(6628)

because the lines are parallel (they never cross)

the equations are in slope-intercept form (y = mx + b)

the lines have the same slope (m), which makes them parallel

they are not the same line, because the y-intercepts (b) are different